Shift Summary for Eve shift, 07.28.99

Jim Graham (jagraham@hep.uchicago.edu)
Thu, 29 Jul 1999 00:07:41 -0500

Shift Summary for Eve shift, 07.28.99

On Shift: JAG, MB

Hours of beam delivered at all intensities: 6.25 / 8.5
Hours of beam delivered at nominal intensity: 6.25 / 8.5
Hours of E832 data (special runs not included): 5.60 / 8.5
Hours of special runs (laser scans, muon runs...): none

Accounting of lost time: MI problem beam problems, pipeline/stream
problems, DDD2 veto.

========================================================================
Runs ended this shift:
Run 14048 - 3.3M events - BEAM_832
Run 14049 - 3.0M events - BEAM_832

Tapes brought to FCC this shift:
QK*185-187, QKB063

========================================================================
Accesses during shift: none

========================================================================
Overview of shift:
Not so smooth. Had various problems in roughly serial order that killed
our efficiency. Had intermittent beam to start, then pipeline problems, a
readout stream problem, a DDD2 veto problem.

========================================================================
Description of Shift:

run spill time comment
------------------------------------------------------------------------
14048 4 15:40 Late shiftee (JAG) arrives
20 16:00 Missing spills due to MI sextapole magnet
42 16:30 Beam returns (4.0E12)
Checked off-spill current draw for DC3Y, it was
relatively low (3.4uA).
50 16:42 Trigger histograms missing, but blank window there.
Determine that there was a 'fix_disp' process running
from an old run. Kill that, kill the mon_disp, and
start a new fix_disp.
56 16:48 Beam gone
17:55 Beam back
115 Good data (beam position) starts
148 Massive ktevana errors. Pipeline init and kfrend
After this point, we lose pipeline quadrant UL, and
no amount of pipeline-initting brings it back. In
consultation w/ Run Manager, we stop the run and
do a fresh start (for fix_disp)
19:18 End run

14049 1 19:23 Start run. Fixes pipeline/stream problem
58 20:43 Run hangs on DDD 2 veto. Try to stop gracefully for
EOFs, but fail to get one on QKK187. New Tapes and
a drt_fresh_start.

21:12 SY Energy conservation
21:23 Beam back

14051 1 21:30 Start run
21:45 Call MCR about SEM ~4.6-4.7 E 12. They can't do
anything because our 'dump' (meson line) lost
it's interlocks, and we are getting all the beam.

00:00 Hand off to Val and Lisa.

========================================================================
Follow up Notes on Previous Shift

========================================================================
Items to Follow up on Next Shift